# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

### v2

@@ Abstract.
We investigate the large deviations principle (which concerns sequences of exponentially small sets) for the isoperimetric problem on product Riemannian manifolds $\mathcal{M}^{n}$ equipped with product probability measures $\nu^{\otimes n}$, where $\mathcal{M}$ is a Riemannian manifold satisfying curvature-dimension bound $\mathrm{CD}(0,\infty)$. When the probability measure $\nu$} admits a finite moment generating function for squared distance, we establish an exact characterization of the large deviations asymptotics for the isoperimetric profile, which shows a precise equivalence between these asymptotic isoperimetric inequalities and nonlinear log-Sobolev inequalities. It is observed that the product of two relative entropy typical sets (or empirically typical sets) forms an asymptotically optimal solution to the isoperimetric problem. The proofs in this paper integrate tools from information theory, optimal transport, and geometric measure theory.
@@ 1. Introduction
In a measure-metric space, the isoperimetric problem concerns determining which subset has the smallest surface area (formally defined via boundary measure) given a fixed measure of the subset. This classic problem remains largely unsolvable in its exact form—explicitly characterizing the isoperimetric profile (the minimal surface area for each fixed measure) is widely open except for specific cases (e.g., standard Gaussian measures, log-concave measures on the real line, or Euclidean spheres with constant density). Instead of pursuing exact solutions, this paper focuses on asymptotic solutions to the isoperimetric problem, particularly in the high-dimensional regime. Specifically, we consider the following setting.
Let $\Omega:=(\mathcal{M},g)$ denote a finite dimensional smooth complete oriented connected Riemannian manifold with induced geodesic distance $d$. Let $\nu=e^{-V}\mathrm{vol}$ with $V\in C^{2}(\mathcal{M})$ be a reference Borel probability measure on $\mathcal{M}$ that is absolutely continuous with respect to the Riemannian volume $\mathrm{vol}$. We consider the $n$-fold product space $\mathcal{M}^{n}$, equipped with the product measure $\nu_{n}:=\nu^{\otimes n}$ and the product distance $d_{n}(\mathbf{x},\mathbf{y})=\sqrt{\sum_{i=1}^{n}d(x_{i},y_{i})^{2}}$, where $\mathbf{x}=(x_{1},x_{2},...,x_{n})$.
@@
where $A$ is a Borel set. Note that there are several other ways in the literature to define the perimeter of a set, but they coincide when the set admits smooth boundary, and thus, lead to the same isoperimetric profile in our setup; see details in [6, pp. 32-33].
Exactly characterizing the isoperimetric profile is in general extremely difficult, and remains widely open except for certain specific cases. For example, when $\nu$ is the standard Gaussian measure, the isoperimetric profile $I_{n}(a)=I_{\mathrm{G}}(a):=\varphi(\Phi^{-1}(a))$ for all $n\geq 1$ with half spaces as isoperimetric minimizers [43, 13]. Here, $\varphi$ and $\Phi$ are respectively the pdf and cdf of the standard Gaussian. It is widely known that $I_{\mathrm{G}}(a)\sim a\sqrt{2\ln\frac{1}{a}}$ as $a\to 0$. That is, the logarithmic isoperimetric constant for the Gaussian measure is
$\lim_{a\to 0}\frac{I_{\mathrm{G}}(a)}{a\sqrt{\ln\frac{1}{a}}}=\sqrt{2}.$
@@
Due to the difficulty of exactly determining the isoperimetric profile especially in the high dimension case, in this paper we aim at investigating the asymptotic behavior, specifically, large deviations asymptotics of the isoperimetric profile. We investigate the asymptotic behavior of the isoperimetric profile when the probability measure of the set decreases exponentially as the dimension increases. That is, we determine the following limits: for $\alpha>0$,
$\underline{\Lambda}(\alpha)$ $:=\liminf_{n\to\infty}\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}\sqrt{n}},$
(1.4) $\overline{\Lambda}(\alpha)$ $:=\limsup_{n\to\infty}\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}\sqrt{n}}.$
We denote them by $\Lambda(\alpha)$ if they are equal.
@@ 1.1.1. Large Deviations Principle
(1.5) $\mathrm{Ric}_{g}+\mathrm{Hess}_{g}V\geq 0\;\;\text{ as 2-tensor fields},$
where $\mathrm{Ric}_{g}$ is the Ricci curvature and $\mathrm{Hess}_{g}V$ is the Hessian of $V$. Since $I_{n}(a)/a$ is non-increasing in $a$ under the assumption (1.5) [32, Proposition 3.1], $\Lambda$ is non-decreasing.
@@
Assume that (1.5) holds. Then, $\Lambda$ is non-decreasing.
Let $P_{2}^{\mathrm{ac}}(\mathcal{M})$ be the set of probability measures which are absolutely continuous with respect to $\nu$ and have finite second-order moments. We also need the concept of nonlinear log-Sobolev inequality which was first introduced by Polyanskiy and Samorodnitsky [39], as well as the concept of typical sets which are fundamental in information theory [15].
@@
Any convex $\theta$ admiting a $\theta$-nonlinear log-Sobolev inequality must satisfy that $\theta(0)=0$ and $\theta$ is non-decreasing. In particular, when $\theta(t)=Kt$ for some constant $K>0$, then the $\theta$-nonlinear log-Sobolev inequality reduces to the classic log-Sobolev inequality. Define
@@
Given $n\in\mathbb{N},\epsilon>0$, the (relative entropy-) typical set of $\mu$ w.r.t. $\nu$ is given by
$A_{n,\epsilon}(\mu\|\nu):=\left\{\mathbf{x}\in\mathcal{M}^{n}:\frac{1}{n}\sum_{i=1}^{n}\ln\frac{{\text{\rm d}}\mu}{{\text{\rm d}}\nu}(x_{i})\geq D(\mu\|\nu)-\epsilon\right\}.$
The quantity $\sum_{i=1}^{n}\ln\frac{{\text{\rm d}}\mu}{{\text{\rm d}}\nu}(x_{i})$ is known as the information density of $\mu^{\otimes n}$ w.r.t. $\nu^{\otimes n}$. So, the one-sided typical set is the half-space in the information density domain. This concept generalizes the half-spaces $\left\{\mathbf{x}\in\mathbb{R}^{n}:\frac{1}{n}\sum_{i=1}^{n}x_{i}\leq b\right\}$ in Euclidean spaces with $b\in\mathbb{R}$, and the former reduces to the latter when $\nu=\mathcal{N}(0,1),\mu=\mathcal{N}(c,1)$. It also generalizes the Euclidean balls $\left\{\mathbf{x}\in\mathbb{R}^{n}:\frac{1}{n}\sum_{i=1}^{n}x_{i}^{2}\leq b\right\}$ with $b>0$, and the former reduces to the latter when $\nu=\mathrm{Lebesgue}$ (which is an infinite measure and not considered in this paper) and $\mu=\mathcal{N}(0,c)$ with $c>0$. Note that half-spaces and Euclidean balls are respectively isoperimetric minimizers for Gaussian measures and Lebesgue measures.
We now establish the large deviations principle for the isoperimetric problem, a result of particular interest as it reveals a precise equivalence between the asymptotic isoperimetric inequality and the nonlinear log-Sobolev inequality. We also need the “full exponential integrability of squared distance”: for some $x_{0}\in\mathcal{M}$ (equivalently for any $x_{0}\in\mathcal{M}$),
(1.8) $\mathbb{E}_{\nu}[e^{\lambda d^{2}(X,x_{0})}]<\infty,\;\forall\lambda\in(0,\infty).$
In other words, the probability measure $\nu$} admits a finite moment generating function for squared distance. This condition is satisfied if $\mathcal{M}$ is a bounded Riemann manifold or if $\mathcal{M}$ is a Euclidean space with $\nu$ having density $\propto e^{-C|x|^{\beta}}$ where $C>0$ and $\beta>2$.
@@
Assume that (1.5) and (1.8) hold. Then, it holds that for any $\alpha>0$,
@@
Moreover, the sequence of sets $A_{\lfloor n\lambda^{*}\rfloor,\epsilon}(\mu_{1}^{*}\|\nu)\times A_{\lceil n(1-\lambda^{*})\rceil,\epsilon}(\mu_{2}^{*}\|\nu),n\in\mathbb{N}$ asymptotically attains the limit $\Lambda(\alpha)$ as $n\to\infty$ first and $\epsilon\to 0$ then, where $(\lambda^{*},\mu_{1}^{*},\mu_{2}^{*})$ is the optimal solution in (1.7).
@@
Although this theorem does not include the Gaussian measure case and the Lebesgue measure case, for these two cases, the one-sided typical sets with $\mu$ set to the extremers (see [39]) respectively reduce to half-spaces and Euclidean balls, consistent with the already known fact that these two kinds of sets are isoperimetric extremers for these two cases. Moreover, by this already known fact, Equation (1.9) still holds for these two cases, suggesting that Theorem 1 might hold for a much larger class of measures.
This theorem can be alternatively formulated as follows: For any convex $\theta:[0,\infty)\to[0,\infty)$, the following two statements are equivalent.
  • $\Lambda(\alpha)\geq\sqrt{\theta(\alpha)}$} holds for all $\alpha>0$.
  • $I(\mu\|\nu)\geq\theta(D(\mu\|\nu)),\;\forall\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M})$.
@@
where $a_{n}=e^{-n\alpha}$ and $I_{\mathrm{G}}$ is the isoperimetric profile for the standard Gaussian measure. That is, as the volume vanishes exponentially, the asymptotic behavior of the isoperimetric profile for any product probability measure resembles the one for the standard Gaussian measure, except that they differ only in a factor determined by the rate of the exponential convergence of the volume.
@@
Theorem 1 reveals a precise equivalence between the asymptotic isoperimetric inequality and the nonlinear log-Sobolev inequality. It is termed the Large Deviations Principle for Isoperimetry for two reasons. First, both this theorem and large deviations theory focus on sets of exponentially small measures. Second, isoperimetric minimizers have been proven to be expressed in terms of one-sided typical sets which concern the remote tails of some sequences. These remote tails themselves are the core object of investigation in large deviations theory.
@@
Showing the equivalence among inequalities from various fields is an interesting topic, since it will unveil the intrinsic relationships among diverse branches of mathematics. However, most of equivalences in the literature are known only in a weak sense. In particular, it was proved by Ledoux [25, 26] on Gaussian spaces or compact Riemannian manifolds and by Bobkov [10] on Euclidean spaces with any log-concave probability measures that for a fixed $n$,
@@
holds for some $K>0$, if and only if the classic log-Sobolev inequality holds with some constant $K^{\prime}>0$. Theorem 1 provides an equivalence in a stronger sense, which is applied in the following to compare the optimal constants in various inequalities.

### v3

@@ Abstract.
The isoperimetric problem is a classic topic in geometric measure theory, yet critical questions regarding the characterization of optimal solutions—even asymptotically optimal ones—remain largely unresolved. In this paper, we investigate the large deviations asymptotics (governing sequences of exponentially small sets) for the isoperimetric problem on the product Riemannian manifold $\mathcal{M}^{n}$ endowed with the product probability measure $\nu^{\otimes n}$, where $\mathcal{M}$ is a Riemannian manifold satisfying the curvature-dimension condition $\mathrm{CD}(0,\infty)$. When the probability measure $\nu$} admits a finite moment generating function for squared distance, we establish an exact characterization of the large deviations asymptotics of the isoperimetric profile, which reveals a precise equivalence between this asymptotic isoperimetric inequality and an established strengthening of log-Sobolev inequality—namely, the nonlinear log-Sobolev inequality. It is observed that conditional typical sets (i.e., $\mathbf{f}$-coordinate half-spaces for some vector $\mathbf{f}$ of functions), a fundamental concept from information theory, form an asymptotically optimal solution to the isoperimetric problem. This class of subsets further yields an upper bound on the isoperimetric profile in the central limit regime (concerning constant-volume sets), expressed as the Gaussian isoperimetric profile times the square root of the spectral gap. Furthermore, we apply our results to establish quantitative relations among the optimal constants in isoperimetric, concentration and transport inequalities, additionally covering isoperimetry for subexponentially and superexponentially small sets. Our results provide a rigorous justification from the perspective of nonlinear log-Sobolev inequalities for why isoperimetric minimizers behave fundamentally differently across spaces with distinct geometric structures. We adopt a new proof idea in this paper which integrates tools from information theory, optimal transport, and geometric measure theory.
@@ 1. Introduction
In a metric-measure space, the isoperimetric problem seeks the subset of fixed measure that minimizes surface area—where surface area is rigorously defined via the boundary measure. This classic problem remains largely unresolved in both its exact and asymptotic formulations, with the explicit characterization of the isoperimetric profile (the minimal surface area corresponding to each fixed measure) and its asymptotic behavior posing a widely open question, save for a handful of specific cases: e.g., the standard Gaussian measure, log-concave measures on the real line, and Euclidean spheres endowed with constant density. In this paper, we study the isoperimetric problem across various asymptotic regimes, with particular emphasis on the large deviations regime that considers sequences of exponentially small sets. We formalize our core analytical setup as follows.
Let $(\mathcal{M},g)$ be a finite-dimensional smooth complete oriented connected Riemannian manifold with induced geodesic distance $d$. Let $\nu=e^{-V}\mathrm{vol}$ with $V\in C^{2}(\mathcal{M})$ be a reference Borel probability measure on $\mathcal{M}$ that is absolutely continuous with respect to the Riemannian volume $\mathrm{vol}$. We consider the $n$-fold product space $\mathcal{M}^{n}$, equipped with the product measure $\nu_{n}:=\nu^{\otimes n}$ and the product distance $d_{n}(\mathbf{x},\mathbf{y})=\sqrt{\sum_{i=1}^{n}d(x_{i},y_{i})^{2}}$, where $\mathbf{x}=(x_{1},x_{2},...,x_{n})$.
@@
where $A$ is a Borel set. Note that there are several other ways in the literature to define the perimeter of a set, but they coincide when the set admits smooth boundary, and thus, lead to the same isoperimetric profile in our setup; see details in [10, 5, pp. 32-33].
Exactly characterizing the isoperimetric profile is in general extremely difficult, and remains widely open except for certain specific cases. For example, when $\nu$ is the standard Gaussian measure, the isoperimetric profile $I_{n}(a)=I_{\mathrm{G}}(a):=\phi(\Phi^{-1}(a))$ for all $n\geq 1$ with half-spaces as isoperimetric minimizers [61, 18]. Here, $\phi$ and $\Phi$ are respectively the pdf and cdf of the standard Gaussian. It is widely known that $I_{\mathrm{G}}(a)\sim a\sqrt{2\ln\frac{1}{a}}$ as $a\to 0$. That is, the asymptotic isoperimetric ratio for the Gaussian measure is
(1.4) $\lim_{a\to 0}\frac{I_{\mathrm{G}}(a)}{a\sqrt{\ln\frac{1}{a}}}=\sqrt{2}.$
@@
As noted at the outset of this paper, our primary aim is to study the asymptotic behavior of the isoperimetric profile, in particular its large deviations asymptotics. We focus on this behavior in the regime where the probability measure of the set decays exponentially with increasing dimension, and to this end we identify the following limits: for $\alpha>0$,
(1.5) $\underline{\Lambda}(\alpha)$ $:=\liminf_{n\to\infty}\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}\sqrt{n}},$
and $\overline{\Lambda}(\alpha)$ which is defined similarly but with $\liminf_{n\to\infty}$ replaced by $\limsup_{n\to\infty}$. We denote them by $\Lambda(\alpha)$ if they are equal.
@@ 1.1.1. Large Deviations Principle
$\mathrm{Ric}_{g}+\mathrm{Hess}_{g}V\geq 0\;\;\text{ as 2-tensor fields},$
where $\mathrm{Ric}_{g}$ is the Ricci curvature and $\mathrm{Hess}_{g}V$ is the Hessian of $V$.
Since $I_{n}(a)/a$ is non-increasing in $a$ under Assumption 1 [45, Proposition 3.1], $\Lambda$ is non-decreasing.
@@
Under Assumption 1, $\Lambda$ is non-decreasing.
Let $P_{2}^{\mathrm{ac}}(\mathcal{M})$ be the set of probability measures which are absolutely continuous with respect to $\nu$ and have finite second-order moments. We also need the concept of nonlinear log-Sobolev inequality and a generalized version of half-spaces.
@@
The nonlinear log-Sobolev inequality, also known as the entropy-energy inequality or the entropy isoperimetric inequality, can be seen as a strengthening of the standard (linear) log-Sobolev inequality since it provides more information about the tradeoff between entropy and Fisher information. This concept was investigated in several important works in the literature, e.g., Costa and Cover [20], F. Y. Wang [66], Bakry, Gentil, and Ledoux [7], as well as Polyanskiy and Samorodnitsky [55].
Any convex $\theta$ that admits a $\theta$-nonlinear log-Sobolev inequality must satisfy that $\theta(0)=0$ and $\theta$ is non-decreasing. In particular, when $\theta(t)=Kt$ for some constant $K>0$, then the $\theta$-nonlinear log-Sobolev inequality reduces to the standard log-Sobolev inequality. Define
@@
The second concept we need is a generalized version of half-spaces which incorporates many important concepts as special cases, e.g., the concept of typical sets from information theory.
@@
Given $n\in\mathbb{N}$ and $\mathbf{f}=(f_{1},f_{2},...,f_{n})$ with each $f_{i}:\mathcal{M}\to\mathbb{R}$, the $\mathbf{f}$-coordinate half-space is defined as
(1.8) $H_{t}(\mathbf{f}):=\Big\{\mathbf{x}\in\mathcal{M}^{n}:\sum_{i=1}^{n}f_{i}(x_{i})\geq t\Big\}$
for a real number $t$. In particular, when $f_{i}=\ln\frac{{\text{\rm d}}\mu_{i}}{{\text{\rm d}}\nu}$ for some probability measures $\mu_{i},i\in[n]$ and $t=\sum_{i=1}^{n}D(\mu_{i}\|\nu)-n\epsilon$ for some $\epsilon>0$, the $\mathbf{f}$-coordinate half-space reduces to the conditional (relative entropy) typical set of $(\mu_{i})_{i\in[n]}$ w.r.t. $\nu$ which is given by
$A_{n,\epsilon}((\mu_{i})_{i\in[n]}\|\nu):=\Big\{\mathbf{x}\in\mathcal{M}^{n}:\sum_{i=1}^{n}\ln\frac{{\text{\rm d}}\mu_{i}}{{\text{\rm d}}\nu}(x_{i})\geq\sum_{i=1}^{n}D(\mu_{i}\|\nu)-n\epsilon\Big\}.$
Remark 2.
The concept of $\mathbf{f}$-coordinate half-space reduces to the Orlicz ball if $f_{i}$ are Young functions (and t>0-t>0). However, we do not pose any restriction on functions $f_{i}$ here (except for measurability), and thus our concept is much more general and incorporates standard half-spaces as special cases.
Remark 3.
Here, $(\mu_{i})_{i\in[n]}$ can be seen as a conditional probability measure $\mu_{X|W}$ with $\mu_{X|W=i}=\mu_{i}$, $\frac{1}{n}\sum_{i=1}^{n}\ln\frac{{\text{\rm d}}\mu_{i}}{{\text{\rm d}}\nu}(x_{i})$ is known as the conditional information density of $\mu_{X|W}$ w.r.t. $\nu$ given $\mu_{W}\sim\mathrm{Unif}[n]$, and its expectation is $D(\mu_{X|W}\|\nu|\mu_{W}):=\frac{1}{n}\sum_{i=1}^{n}D(\mu_{i}\|\nu)$ which is known as the conditional relative entropy. That is why it is called the conditional typical set. In addition, the standard definition of the conditional relative entropy typical set is $\big\{\mathbf{x}\in\mathcal{M}^{n}:\big|\sum_{i=1}^{n}\ln\frac{{\text{\rm d}}\mu_{i}}{{\text{\rm d}}\nu}(x_{i})-\sum_{i=1}^{n}D(\mu_{i}\|\nu)\big|\leq n\epsilon\big\}$ [21, (11.206)]. Our definition here is slightly different from the standard one, and has benefit that it incorporates the isoperimetric minimizers for Gaussian and Lebesgue measures as special cases, which is explained below.
The concept of $\mathbf{f}$-coordinate half-space generalizes the standard half-spaces $\{\mathbf{x}\in\mathbb{R}^{n}:\sum_{i=1}^{n}w_{i}x_{i}\leq t\}$ in Euclidean spaces, and the former reduces to the latter when $f_{i}(x_{i})=-w_{i}x_{i}$ and $t$ is rechosen, or essentially equivalently, set $\nu=\mathcal{N}(0,1),\mu_{i}=\mathcal{N}(w_{i},1)$ in the conditional typical set. It also generalizes the Euclidean balls $\left\{\mathbf{x}\in\mathbb{R}^{n}:\sum_{i=1}^{n}x_{i}^{2}\leq t\right\}$ with $t>0$, and the former reduces to the latter when $f_{i}(x_{i})=-x_{i}^{2}$ and $t$ is rechosen, or essentially equivalently, set $\nu=\mathrm{Lebesgue},\mu=\mathcal{N}(0,c)$ with $c>0$ in the conditional typical set (although the Lebesgue measure is an infinite measure and not considered in this paper). Note that standard half-spaces and Euclidean balls are respectively isoperimetric minimizers for Gaussian measures and Lebesgue measures.
We now establish the large deviations principle for the isoperimetric problem, a result of particular interest as it reveals a precise equivalence between the asymptotic isoperimetric inequality and the nonlinear log-Sobolev inequality. We also need the following assumption.
Assumption 2 (Full Exponential Integrability of Squared Distance).
For some $x_{0}\in\mathcal{M}$ (equivalently for any $x_{0}\in\mathcal{M}$),
$\mathbb{E}_{\nu}[e^{\lambda d^{2}(X,x_{0})}]<\infty,\;\forall\lambda\in(0,\infty).$
In other words, the probability measure $\nu$} admits a finite moment generating function for squared distance.
This assumption is satisfied if $\mathcal{M}$ is a bounded Riemann manifold or if $\mathcal{M}$ is a Euclidean space with $\nu$ having density $\propto e^{-C|x|^{\beta}}$ where $C>0$ and $\beta>2$.
@@
Under Assumptions 1 and 2, it holds that for any $\alpha>0$,
@@
Moreover, a sequence of conditional typical sets $A_{n,\epsilon}((\mu_{1}^{*},...,\mu_{1}^{*},\mu_{2}^{*},...,\mu_{2}^{*})\|\nu)$ with $\mu_{1}^{*},\mu_{2}^{*}$ respectively appearing $\lfloor n\lambda^{*}\rfloor$ and $n-\lfloor n\lambda^{*}\rfloor$ times asymptotically attains the limit $\Lambda(\alpha)$ as $n\to\infty$ first and $\epsilon\to 0$ then, where $(\lambda^{*},\mu_{1}^{*},\mu_{2}^{*})$ is the optimal solution in (1.7). In addition, the inequality $\overline{\Lambda}(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}$} holds even without Assumption 2.
@@
Although this theorem does not include the Gaussian measure case and the Lebesgue measure case, for these two cases, the conditional typical sets with $\mu$ set to the extremizers (see [55]) respectively reduce to standard half-spaces and Euclidean balls, consistent with the already known fact that these two kinds of sets are isoperimetric extremizers for these two cases. Moreover, by this already known fact, Equation (1.9) still holds for these two cases, suggesting that Theorem 1 might hold for a much larger class of measures.
This theorem can be alternatively formulated as follows: For any convex $\theta:[0,\infty)\to[0,\infty)$,
${\displaystyle\Lambda(\alpha)\geq\sqrt{\theta(\alpha)}},\;\forall\alpha>0\quad\Longleftrightarrow\quad I(\mu\|\nu)\geq\theta(D(\mu\|\nu)),\;\forall\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M}).$
@@
where $a_{n}=e^{-n\alpha}$ and $I_{\mathrm{G}}$ is the isoperimetric profile for the standard Gaussian measure. That is, as the volume vanishes exponentially, the asymptotic behavior of the isoperimetric profile for any product probability measure resembles the one for the standard Gaussian measure, but with different factors that are determined by the exponential convergence rate of the volume. One prototypical case occurs when $\Theta(\alpha)$ grows linearly in $\alpha$—for instance, $\Theta(\alpha)=2\alpha$ for Gaussian measures—implying that $I_{n}(a)$ behaves as $C\cdot a\sqrt{\ln\frac{1}{a}}$, with standard half-spaces likely serving as approximate extremizers for small volumes. A second prototypical case is when $\Theta(\alpha)$ grows exponentially in $\alpha$, as illustrated in the subsequent example, which implies that $I_{n}(a)$ behaves as $C\cdot\sqrt{n}\,a^{1-1/n}$, with geodesic balls acting as approximate extremizers for small volumes. Theorem 1 provides a rigorous justification from the perspective of nonlinear log-Sobolev inequalities for why isoperimetric minimizers of small volumes behave fundamentally differently across spaces with distinct geometric structures.
Example 1 (Exponential Growth of $\Theta$).
Let $D(t)=D(\mu_{t}\|\nu)$, $I(t)=I(\mu_{t}\|\nu)$, and $N(t)=\exp(-2D(\mu_{t}\|\nu))$ respectively be the relative entropy, Fisher information, and entropy power along the (generalized) heat flow $\partial_{t}\rho=L\rho$ with $L=\Delta-\nabla V\cdot\nabla$ and $\mu_{t}=\rho_{t}\nu$. It is easy to see that $\Theta(\alpha)$ grows exponentially as $\alpha\to\infty$, as long as $N(t)$ is concave in the time $t$ and $\Theta(\alpha_{0})>0$ for some $\alpha_{0}>0$. This is because, in this case, $D(t)$ continuously decays to zero as $t\to\infty$ and $N^{\prime}(t)=2N(t)I(t)$ is nonincreasing, which implies that for any $\mu_{0}$ satisfying $D(0)\geq\alpha_{0}$, we always have $N(0)I(0)\geq N(t_{0})I(t_{0})\geq\exp(-2\alpha_{0})\Theta(\alpha_{0})$ with $t_{0}$ being the solution to $D(t_{0})=\alpha_{0}$. Therefore, $I(0)\geq\Theta(\alpha_{0})\exp(2(D(0)-\alpha_{0}))$ for any $\mu_{0}$ satisfying $D(0)\geq\alpha_{0}$, i.e., exponential growth of $\Theta$: $\Theta(\alpha)\geq\Theta(\alpha_{0})\exp(2(\alpha-\alpha_{0}))$ for all $\alpha\geq\alpha_{0}$. The concavity of the entropy power holds under the $\mathrm{CD}(0,m)$-condition with $m\in[n,\infty)$, as shown by S. Li and X. Li [42, Theorem 2.2]. The condition that $\Theta(\alpha_{0})>0$ for some $\alpha_{0}>0$ is guaranteed as long as a standard log-Sobolev inequality (or a Poincaré inequality) is admitted. These two conditions are satisfied by the (normalized) unweighted measure on a bounded manifold with nonnegative Ricci curvature, which hence admits exponential growth of $\Theta$ and also exponential growth of $\Lambda$ (since Assumptions 1 and 2 are satisfied in this setting and Theorem 1 is valid). See other sufficient conditions to ensure exponential growth of $\Theta$, e.g. , in [7, Theorem 6.8.1][42, Theorem 2.5].
We next consider a simple but nontrivial case—the uniform measure on hypercube.
Example 2 (Hypercube).
Consider the hypercube $(0,1)^{n}$ with Euclidean metric and uniform measure. The standard log-Sobolev inequality and the Poincaré inequality with constants $K_{LS}=K_{P}=\pi^{2}$ ($K_{LS},K_{P}$ defined later in (1.12) and (1.18)) respectively hold in this setting [27], which implies $\lim_{\alpha\to 0}\frac{\breve{\Theta}(\alpha)}{2\alpha}=\lim_{\alpha\to 0}\frac{\Theta(\alpha)}{2\alpha}=\pi^{2}$. Moreover, the $\mathrm{CD}(0,m)$-condition holds obviously. Thus, as mentioned in Example 1, we have exponential growth of $\Theta$, i.e., $\Theta(\alpha)\geq 2\pi^{2}\alpha_{0}\cdot e^{2(\alpha-\alpha_{0})}\geq\frac{\pi^{2}}{e}\cdot e^{2\alpha}$ (with $\alpha_{0}$ chosen $1/2$), which further implies exponential growth of $\Lambda$, i.e., $\Lambda(\alpha)\geq\sqrt{\frac{\pi^{2}}{e}}\cdot e^{\alpha}$. This lower bound is sharp in the exponent by verifying balls centered at the vertices of the hypercube, since for these balls with volume $e^{-n\alpha}$, their surface areas are $(1+o(1))\sqrt{\frac{\pi en}{2}}\cdot e^{\alpha(1-n)}$ as $n\to\infty$, giving $\Lambda(\alpha)\leq\sqrt{\frac{\pi e}{2}}\cdot e^{\alpha}$. In fact, it was already known [58, 29] that given each $n$, for all small volumes, balls centered at the vertices are isoperimetric minimizers. However, it is unclear whether “small volume” here is large enough to include “exponentially small volume”. We conjecture the following precise relation holds: $\Lambda(\alpha)\sim\sqrt{\frac{\pi e}{2}}\cdot e^{\alpha}$, or equivalently, $\breve{\Theta}(\alpha)\sim\frac{\pi e}{2}\cdot e^{2\alpha}$ as $\alpha\to\infty$, suggesting that half-normal distributions truncated on $(0,1)$ with sufficiently small variances approximately saturate nonlinear log-Sobolev inequalities for large $\alpha$. The factor $\frac{\pi^{2}}{e}$ obtained above is about $0.85\cdot\frac{\pi e}{2}$.
@@
Theorem 1 reveals a precise equivalence between the asymptotic isoperimetric inequality and the nonlinear log-Sobolev inequality. It is termed the large deviations principle for isoperimetry for two reasons. First, both this theorem and large deviations theory focus on sets of exponentially small measures. Second, isoperimetric minimizers have been proven to be expressed in terms of $\mathbf{f}$-coordinate half-spaces (or conditional typical sets) which concern the remote tails of a sum. These remote tails themselves are the core object of investigation in large deviations theory.
@@
Showing the equivalence among inequalities from various fields is an interesting topic, since it will unveil the intrinsic relationships among diverse branches of mathematics. However, most of equivalences in the literature on this topic are known only in a weak sense. In particular, it was proved by Ledoux [36, 37] on Gaussian spaces or compact Riemannian manifolds and by Bobkov [14] on Euclidean spaces with any log-concave probability measures that for a fixed $n$,
@@
holds for some $K>0$, if and only if the standard log-Sobolev inequality holds with some constant $K^{\prime}>0$. Theorem 1 provides an equivalence in a stronger sense for product spaces, which showing a precise equivalence between the isoperimetric inequality and (nonlinear) log-Sobolev inequality. Our equivalence will be applied in the following to compare the optimal constants in various inequalities.